OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4

OpenStudy (yrelhan4):

opening the bracket on the LHS -2x - 6 = -2x -2 -4 Now you just need to solve it.

OpenStudy (yrelhan4):

Eh, wait. there is some problem with this question. I'd like you to recheck the question.

OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (yrelhan4):

Ok now. when you solve this statement, you get 1=1. This means that the statement is true for all x.

OpenStudy (anonymous):

x=1

OpenStudy (adilalvi):

-2x-6=-2x-2-4 -2x+2x-6+6=0 0=0 answer

OpenStudy (yrelhan4):

I am sorry, but x=1 is not the answer. This statement is true for all x, as @adilalvi solved and explained by me earlier.